OpenStudy (anonymous):

3x + 4y =8 -2x + 5y = 3 i got no solution. am i right or wrong? if wrong what is the correct answer?

OpenStudy (anonymous):

system of equations problem ~

OpenStudy (kamille):

well, i do get a solution

OpenStudy (anonymous):

what'd you get?

OpenStudy (anonymous):

3x + 4y =8 -2x + 5y = 3 multiply the top equation by 2 multiply the bottom equation by 3 this allows the x values to cancel out since the top is +6x and the bottom is -6x 2(3x+4y=8) 3(-2x+5y=3) 6x+8y=16 -6x+15y=9 add vertically. 8y+15y=16+9 23y=25 y = 25/23 plug back into either equation to find x.

OpenStudy (kamille):

i did it the same way as tomo:)

OpenStudy (anonymous):

i did it again and got 1.2 and 1.1 for the answers. o.O